OpenStudy (anonymous):

Solve using S=r(theta) Radius Central Angle Arc length ? pi/3 3/2 m

OpenStudy (anonymous):

its 3/2=pi/3(r) i dont know how to find r

OpenStudy (anonymous):

@robtobey please please please help me

OpenStudy (anonymous):

i obviously know that i have to get r by itself but i dont know how cause the pi is confusing me.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so you want to solve \[\Large \frac{3}{2} = \frac{\pi}{3}r\] for r?

jimthompson5910 (jim_thompson5910):

if so, multiply both sides by the reciprocal of pi/3 \[\Large \Large \frac{3}{2} = \frac{\pi}{3}r\] \[\Large \Large \color{red}{\frac{3}{\pi}}*\frac{3}{2} = \color{red}{\frac{3}{\pi}}*\frac{\pi}{3}r\]

OpenStudy (anonymous):

i just keep getting a decimal.

OpenStudy (anonymous):

oooooooooo..... 9/2pi

jimthompson5910 (jim_thompson5910):

yeah I'd leave it as a fraction and leave it in terms of pi

OpenStudy (anonymous):

\[\pi d=\text{circumfrence} \]\[\pi 2 r=\text{circumfrence} \]\[\frac{\pi 2 r}{6}=\frac{\text{circumfrence}}{6} \]\[\frac{\pi 2 r}{6}=\frac{3}{2} \]Solve for r.

jimthompson5910 (jim_thompson5910):

\[\Large r = \frac{9}{2\pi}\]

OpenStudy (anonymous):

That is what Mathematica calculated.